Write an equation for the line containing (-6, -6) and (-3, 1)?
Answer by Alan3354(69443)

Here are 2 examples using 2 different methods.
Determinants is much simpler, imo.
-----------------
Write the equation of the line AB that passes through the following points:
A(-2,4) B(5,6)
Find the slope, m = diffy/diffx
m = (4-6)/(-2-5) = -2/-7
m = 2/7
--------
Use y = mx + b and either point to find b
6 = (2/7)*5 + b
b = 6 - 10/7
b = 32/7
y = mx + b
y = (2/7)x + 32/7 Slope-intercept form
2x - 7y = -32 Standard form
--------------------
and A(-3,-4) B(2,8)
Do it like the 1st one, or
---------------------------------
Using determinants
---------------
|+x +y +1|
|-3 -4 +1| = 0
|+2 +8 +1|
--------------
x(-4-8) -y(-3-2) + (-24+8) = 0
-12x + 5y = 16
12x - 5y = -16